The solubility of AgCl at 25 C is 1.3 x 10 -5 moles per liter. Express this as a solubility in grams per 100mL. Show the calculations here.
1
Expert's answer
2020-11-09T13:51:03-0500

Solubility in mol/mL ="\frac{1.3×10^-5mol }{L} × \frac{143.5g AgCl}{mol} ×\frac{1L}{10×100mL}"

= 1.87×10-4 mol/ 100mL